OpenStudy (chrisplusian):

can someone help with an integral problem involving partial fraction decomposition please!!!!!!

OpenStudy (chrisplusian):

\[\int\limits \frac{ x+2 }{ (x^2)-4x } dx\] I got the solution using another method, checked with the solution guide and it worked, problem is I have a test on this and I emailed the professor and she said that if the directions said to use partial fraction that no other method was acceptable. when I try to do partial faction decomposition I get an incorrect answer:\[\int\limits \frac{ x+2 }{ x(x-4) } dx\] then I got \[\int\limits \frac{ x+2 }{ x(x-4) } dx =\frac{ a }{ x }+\frac{ b }{ x-4 }\] when I solve I end up with \[\int\limits \frac{ 1 }{ x } +\frac{ 0 }{ x-4 } dx\] which is just a natural log of x plus c as a result....... not the correct answer

zepdrix (zepdrix):

A zero for b? Hmm let's see what went wrong.

zepdrix (zepdrix):

 Setup look ok so far? :)